# Generative Text Steganography with Large Language Model

### v1

@@ Abstract.
Recent advances in large language models (LLMs) have blurred the boundary of high-quality text generation between humans and machines, which is favorable for generative text steganography. While, current advanced steganographic mapping is not suitable for LLMs since most users are restricted to accessing only the black-box API or user interface of the LLMs, thereby lacking access to the training vocabulary and its sampling probabilities. In this paper, we explore a black-box generative text steganographic method based on the user interfaces of large language models, which is called LLM-Stega. The main goal of LLM-Stega is that the secure covert communication between Alice (sender) and Bob (receiver) is conducted by using the user interfaces of LLMs. Specifically, We first construct a keyword set and design a new encrypted steganographic mapping to embed secret messages. Furthermore, to guarantee accurate extraction of secret messages and rich semantics of generated stego texts, an optimization mechanism based on reject sampling is proposed. Comprehensive experiments demonstrate that the proposed LLM-Stega outperforms current state-of-the-art methods.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXprice: isbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Human and societal aspects of security and privacyccs: Computing methodologies Natural language generation
@@ 1. Introduction
Steganography is the science and art of embedding secret information into cover media, aiming to covertly transmit secret information through public channels.(Anderson and Petitcolas, 1998; Provos and Honeyman, 2003; Cox et al., 2007). The principle of steganography can be illustrated by Simmons’ ”prisoner problem”(Simmons, 1984): There are two prisoners Alice (sender) and Bob (receiver) in jail who are trying to hatch an escape plan. The only way they can communicate is carefully censored by warden Eve (steganalyzer). Therefore, they must find some way to embed the secret messages into a normal digital carrier (recalled cover) to obtain an “innocent-looking” medium with secret messages (called stego). According to the types of digital carrier, steganography can be divided into image steganography (Lu et al., 2021; Su et al., 2024; Yu et al., 2024; Yao et al., 2024), video steganography (Mou et al., 2023; Meng et al., 2024; Mohamed et al., 2024; Jahromi et al., 2024), audio steganography (Jiang et al., 2020; Wu et al., 2020; AlSabhany et al., 2020) and text steganography (Yang et al., 2018; Ziegler et al., 2019; Dai and Cai, 2019; Peng et al., 2023; Zhang et al., 2021; Ding et al., 2023). Recently, text steganography has become a popular research topic in the information security field since text has high universality and robustness when transmitted through public channels.
@@
In order to address the above questions and fully leverage the generation ability of LLMs, as shown in Figure  1(b), we explore a black-box generative text steganography based on the UIs (user interfaces) of large language models. The contributions of this paper as the following:
@@ 3.1. Keyword Set Construction
Notably, randomly selecting keywords from the four subsets appear with unclear logic and semantic blur, which brings difficulties to the satisfied stego text generation. In order to mitigate the question, the LLM, induced by an evaluation prompt, is used to evaluate the superiority of the randomly selected keywords and optimize their probabilities. Compared with existing white-box methods, using the constructed keyword set to encode secret messages has three significant advantages: (1) the keyword set is separated from the generation process of LLM, users do not require accessing the sampling distribution of LLMs. Meanwhile, the steganographic behavior does not change the sampling distribution of LLM, which improves the security of the generated stego texts. (2) since the sampling probabilities of the keywords are optimized by using the potential knowledge of LLM, the generated stego text based on these keywords can achieve high text quality. (3) in an ideal situation, the keyword set can encode massive secret messages as long as the set is large enough.
@@ 3.2. Encrypted Steganographic Mapping
In practice, the number of commonly used keywords is limited and their sampling distribution does not obey uniform distribution. If these subsets are directly used to encode the secret messages, the embedding capacity and security of the generated stego could not be satisfied. To resolve these problems, we first augment the keyword set. Specifically, we perform repeated sampling to expand the keyword set according to the optimized sampling probabilities. Then, the location indices of keywords in the augmentation set are directly used to encode secret messages. In this part, for a good trade-off between time cost and embedding capacity, we expand the capacity of three subsets (subject, predicate, and object) from 16 to $2^{18}$ words and that of the emotion subset from 3 to $2^{10}$ which can encode $3*18+10=64$ bits secret messages.
In the augmentation keyword set, the more common keyword has a higher sampling probability, encoding more secret messages. The augmentation strategy not only enhances the embedding capacity but also maintains the real sampling probabilities of the keywords encoded secret messages.
Moreover, we notice that the location indices of the keywords are fixed, which leads to the potential exposure risk of the steganographic behavior. To further improve the security of steganographic mapping, an encryption strategy is designed. Concretely, we use a One-Time Password mechanism to implement the encryption. XOR operation is performed using the number of the keyword repetitions and the release time of stego text on online social networks (OSNs)., which is formulated as follows:
@@
Compared with existing steganographic mappings, the proposed encrypted steganographic mapping does not destroy the sampling distribution of LLM in the generation process since the steganographic mapping is independent of the generation process of LLM. Because of this trait, the proposed mapping applies to various LLMs with user interfaces. Due to the page limitation, in this paper, we only leverage the UIs of GPT-4 to evaluate the superiority of the proposed encrypted steganographic mapping.
@@ 3.3. Steganographic Text Generation and Secret Message Extraction
The main goal of the proposed LLM-Stega is that using the user interfaces of LLMs implements steganographic text generation and secret message extraction. After selecting keywords by using the proposed encrypted steganographic mapping, we use an embedding prompt to induce LLMs to generate stego texts. Unlike an explicit extraction algorithm, the extraction based on the LLM cannot ensure a perfectly accurate extraction. Thus, we proposed a feedback optimization mechanism based on reject sampling.
In the generation process, under the guidance of an designed extraction prompt, LLM attempts to output the keywords encoded in the secret messages. If the output keywords have errors, the LLM could return the main reason for these errors and optimize the embedding, generation, and extraction prompts until there are no extraction errors. The feedback optimization mechanism based on reject sampling is shown in Algorithm.1 in detail. In our experiments, making two rejecting sampling ensure the accurate extraction of secret messages embedded into each stego sentence. It is noteworthy that there is a generation prompt behind the reject sampling module shown in Figure. 2. The main reason is that during iterative optimization, the generated stego texts tend to be simple sentences merely containing keywords, so as to ensure accurate extraction. While, the sentence diversity and semantic richness of the generated stego texts are diminished, resulting in secure risk. Therefore, we design and optimize the embedding and generation prompts to ensure rich semantic and accurate extraction, respectively. The optimization mechanism based on reject sampling is shown in Algorithm.1 in detail.
After Bob gains the generated stego texts from OSNs, he (or she) leverages the extraction prompt to extract keywords of the generated stego texts using the UIs of LLM. Finally, according to the shared side information, containing the keyword set, private-key, One-Time-Password mechanism, and the release times of stego texts on OSNs, Bob perfectly decodes the secret messages.
@@ 4. Experiment
In this section, we evaluate the performance of LLM-Stega in terms of text quality, embedding capacity, anti-steganalysis ability, and human evaluation. The details of the experimental setup and result analysis are described in the following sections.
@@ 4.1. Experimental Setup
(1) LLM and Theme information selection. In the experiments, we select the UI of GPT-4 to implement block-box generative text steganography since the GPT-4 is an advanced and widely used LLM. Due to the vivid topics and popularity, the Entertainment News is selected as the theme information to constrain the semantics and context of the generated stego texts.
(2) Baselines. Three advanced generative steganographic methods are rebuilt, including Arithmetic (Ziegler et al., 2019), ADG (Zhang et al., 2021), and Discop (Ding et al., 2023). Due our computation and memory limitation, we use the LSTM model trained on News(Thompson, 2017) data as an off-the-shelf language model. For the Arithmetic-based algorithm, we choose the steganography text generated under the different embedding capacities of 1.39 bit per word (bpw) (AC-2) and 3.99 bpw (AC-6) to compare.
@@ 4.2. Metrics
(5) Human Evaluation. For the human evaluation, we focus on evaluating three key aspects of the generated texts: fluency, coherence, and relevance. These criteria are chosen to reflect the intra-sentence linguistic quality, inter-sentence relatedness and causal dependency, and consistency of the generated texts, respectively. We construct a dataset comprising 100 instances of steganographic and non-steganographic texts generated by each algorithm. Three qualified NLP researchers, proficient in English and with a solid understanding of steganography, are employed as annotators. Each annotator independently rates the texts on a five-point scale (ranging from ’very poor’ to ’very good’) for fluency, coherence, and relevance. The higher evaluation score denotes the better generated stego texts.
@@ 4.3. Results and Analysis
LLM-Stega 13.333 5.93 bpw
Table 1 demonstrates the results of text quality of the generated stego texts. In the exploration of literature (Yang et al., 2020b), it is found that texts on public social networks are written by people of different ages and backgrounds in different ways of expression. This leads to the fact that most of the human-written sentences may not obey the optimal language model and form a large variance. From the security, if the generated stego texts have more similar PPL values with normal sentences, their security is higher. In addition, the higher similarity represents the more similarity between cover and stego sentences, namely, higher security. In our experiments, the average PPL of the normal sentence in the News dataset is 185.64. The “AC-2” can achieve the most similar PPL value and the lowest SS. In addition, the “Discop” algorithm gains the best SS and the maximum gap of PPL. While the proposed LLM-Stega can make the best trade-off between PPL and SS.
Table 3. Generated steganographic examples of the proposed LLM-Stega.
Key-words Stego-text
dancer-announce-song-positive Acclaimed dancer, Luna Moves, unassumingly announced her new song, which intriguingly entwines rhythmic beats with her signature performative flair.
band-praise-song-neutral The band’s praise of their latest song was received with neutrality, a calm before the potential storm of fan reactions.
band-released-exhibition-positive The beloved band joyfully released an immersive exhibition, transporting fans through their rich musical journey, sparking waves of nostalgic happiness across the community.
game-released-episode-positive The innovative game developers ecstatically released a gripping new episode, thrilling gamers and inviting them into captivating new worlds.
singer-performed-episode-positive With a voice that touched souls, the singer performed a memorable episode on the live show, casting a warm, positive light across the enchanted audience.
@@
(4) Human Evaluation. In order to further evaluate the effectiveness of the proposed LLM-Stega, we implement the human evaluation experiment. The details of evaluation scores are illustrated in Figure. 4. The experimental results indicate that the LLM-Stega outperformed the other three algorithms. This suggests that LLM-Stega is more effective in generating steganographic texts that are contextually relevant and less detectable. The results demonstrate LLM-Stega’s superior performance in generating steganographic texts that are fluent, coherent, and relevant. This finding is significant in the context of text steganography, highlighting the potential of LLM-Stega in applications where the imperceptibility of the embedded message is crucial. Besides, we also illustrate some examples of stego text generated by LLM-Stega in Table 3 for qualitative analysis. We find that the stego text is fluent enough, with correct grammar and coherent semantics.
@@ 4.4. Ablation Study
The experimental results are shown in Table 4. We notice that using ”Initial optimization” increases the PPL value and decreases BERT scores, namely, the text quality is reduced. This is because the main goal of “Initial optimization” is to ensure accurately extracting secret messages. In subsequent optimization phases, we aim to improve text quality on the condition that the secret information is extracted accurately, achieving high imperceptibility for the stego-text. Thus, both the PPL and the reject rate have decreased. Besides, the single instance of rejection sampling rate is due to the tuned prompt’s ability to produce text that allows for 100% extraction of the secret information in just one attempt. The experimental results show that the “further optimization” process meets expectations. After ”deep iterations”, the probability of rejection sampling drops to its lowest, while the quality of the text is enhanced to its highest. The experimental results demonstrated that the proposed prompt optimization based on reject sampling significantly improved the quality of the generated stego texts.

### v2

@@ Abstract.
Recent advances in large language models (LLMs) have blurred the boundary of high-quality text generation between humans and machines, which is favorable for generative text steganography. Currently, advanced steganographic mapping is not suitable for LLMs since most users are restricted to accessing only the black-box API or user interface of the LLMs, thereby lacking access to the training vocabulary and its sampling probabilities. In this paper, we explore a black-box generative text steganographic method based on the user interfaces of large language models, which is called LLM-Stega. The main goal of LLM-Stega is to ensure secure covert communication between Alice (sender) and Bob (receiver) by using the user interfaces of LLMs. Specifically, We first construct a keyword set and design a new encrypted steganographic mapping to embed secret messages. Furthermore, an optimization mechanism based on reject sampling is proposed to guarantee accurate extraction of secret messages and rich semantics of generated stego texts. Comprehensive experiments demonstrate that the proposed LLM-Stega outperforms current state-of-the-art methods.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28–November 1, 2024; Melbourne, VIC, Australia.booktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28–November 1, 2024, Melbourne, VIC, Australiaisbn: 979-8-4007-0686-8/24/10doi: 10.1145/3664647.3680562ccs: Security and privacy Human and societal aspects of security and privacyccs: Computing methodologies Natural language generation
@@ 1. Introduction
Steganography is the science and art of embedding secret information into cover media, aiming to covertly transmit secret information through public channels (Anderson and Petitcolas, 1998; Provos and Honeyman, 2003; Cox et al., 2007). The principle of steganography can be illustrated by Simmons’ ”prisoner problem”(Simmons, 1984): There are two prisoners Alice (sender) and Bob (receiver) in jail who are trying to hatch an escape plan. The only way they can communicate is carefully censored by warden Eve (steganalyzer). Therefore, they must find some way to embed the secret messages into a normal digital carrier (called cover) to obtain an “innocent-looking” medium with secret messages (called stego). According to the types of digital carrier, steganography can be divided into image steganography (Lu et al., 2021; Su et al., 2024; Yu et al., 2024; Yao et al., 2024), video steganography (Mou et al., 2023; Meng et al., 2024; Mohamed et al., 2024; Jahromi et al., 2024), audio steganography (Jiang et al., 2020; Wu et al., 2020; AlSabhany et al., 2020) and text steganography (Yang et al., 2018; Ziegler et al., 2019; Dai and Cai, 2019; Peng et al., 2023; Zhang et al., 2021; Ding et al., 2023). Recently, text steganography has become a popular research topic in the information security field since text has high universality and robustness when transmitted through public channels.
@@
In order to address the above questions and fully leverage the generation ability of LLMs, as shown in Figure  1(b), we explore a black-box generative text steganography based on the user interfaces (UIs) of large language models. The contributions of this paper as the following:
@@ 3.1. Keyword Set Construction
Notably, randomly selecting keywords from the four subsets appear with unclear logic and semantic blur, which brings difficulties to the satisfied stego text generation. In order to mitigate the question, the LLM, induced by an evaluation prompt, is used to evaluate the superiority of the randomly selected keywords and optimize their probabilities. Compared with existing white-box methods, using the constructed keyword set to encode secret messages has three significant advantages: (1) the keyword set is separated from the generation process of LLM, users do not require accessing the sampling distribution of LLMs. Meanwhile, the steganographic behavior does not change the sampling distribution of LLM, which improves the security of the generated stego texts. (2) since the keywords and sampling their probabilities are optimized by using the potential knowledge of LLM, the generated stego text based on these keywords can achieve high text quality. (3) in an ideal situation, the keyword set can encode massive secret messages as long as the set is large enough.
@@ 3.2. Encrypted Steganographic Mapping
In practice, the number of commonly used keywords is limited and their sampling distribution does not obey uniform distribution. If these subsets are directly used to encode the secret messages, the embedding capacity and security of the generated stego could not be satisfied. To resolve these problems, we first augment the keyword set. Specifically, we perform repeated sampling to expand the keyword set according to the optimized sampling probabilities. Then, the location indices of keywords in the augmentation set are directly used to encode secret messages. In this part, for a good trade-off between time cost and embedding capacity, we expand the capacity of three subsets (subject, predicate, and object) from 16 to $2^{18}$ words and that of the emotion subset from 3 to $2^{10}$ which can encode $3*18+10=64$ bits secret messages. In the augmentation keyword set, the more common keyword has a higher sampling probability, encoding more secret messages. The augmentation strategy not only enhances the embedding capacity but also maintains the real sampling probabilities of the keywords encoded secret messages.
We notice that the location indices of the keywords are fixed, which leads to the potential exposure risk of the steganographic behavior. To further improve the security of steganographic mapping, an encryption strategy is designed. Concretely, we use a One-Time Password mechanism to implement the encryption. XOR operation is performed using the number of the keyword repetitions and the release time of stego text on online social networks (OSNs), which is formulated as follows:
@@
Compared with existing steganographic mappings, the proposed encrypted steganographic mapping does not destroy the sampling distribution of LLM in the generation process since the steganographic mapping is independent of the LLM generation process. Because of this trait, the proposed mapping applies to various LLMs with user interfaces. Due to the page limitation, in this paper, we only leverage the UIs of GPT-4 to evaluate the superiority of the proposed encrypted steganographic mapping.
@@ 3.3. Steganographic Text Generation and Secret Message Extraction
The main goal of the proposed LLM-Stega is that using the user interfaces of LLMs implements steganographic text generation and secret message extraction. After selecting keywords by using the proposed encrypted steganographic mapping, we use an embedding prompt to induce LLMs to generate stego texts. Unlike an explicit extraction algorithm, the extraction relied on the LLM cannot ensure a perfectly accurate extraction. Thus, we proposed a feedback optimization mechanism based on reject sampling.
In the generation process, under the guidance of a designed extraction prompt, LLM attempts to output the keywords encoded in the secret messages. If the output keywords have errors, the LLM could return the main reason for these errors and optimize the embedding, generation, and extraction prompts until there are no extraction errors. The feedback optimization mechanism based on reject sampling is shown in Algorithm.1 in detail. In our experiments, making two rejecting samples ensures the accurate extraction of secret messages embedded into each stego sentence. It is noteworthy that there is a generation prompt behind the reject sampling module shown in Figure. 2. The main reason is that during iterative optimization, the generated stego texts tend to be simple sentences merely containing keywords, so as to ensure accurate extraction. While, the sentence diversity and semantic richness of the generated stego texts are diminished, resulting in secure risk. Therefore, we design and optimize the embedding and generation prompts to ensure rich semantic and accurate extraction, respectively.
After Bob gains the generated stego texts from OSNs, he (or she) leverages the extraction prompt to extract keywords from the generated stego texts using the LLM UIs. Finally, according to the shared side information, containing the keyword set, private-key, One-Time-Password mechanism, and the release times of stego texts on OSNs, Bob perfectly decodes the secret messages.
@@ 4. Experiment
In this section, we evaluate the performance of LLM-Stega in terms of text quality, embedding capacity, anti-steganalysis ability, statistical imperceptibility and human evaluation. The details of the experimental setup and result analysis are described in the following sections.
@@ 4.1. Experimental Setup
[…]
(2) Baselines. Three advanced generative steganographic methods are rebuilt, including Arithmetic (Ziegler et al., 2019), ADG (Zhang et al., 2021), and Discop (Ding et al., 2023). Due to our computation and memory limitations, we use the LSTM model trained on News(Thompson, 2017) data as an off-the-shelf language model. However, to demonstrate that the advantages of our proposed method are not solely attributed to the use of LLMs, we also implemented a version using Discop with ${}_{\text{LLaMA2}}$) as the language generation model, testing whether LLMs without fine-tuning improves the generation of steganographic texts. For the Arithmetic-based algorithm, we choose the steganography text generated under the different embedding capacities of 1.39 bit per word (bpw) (AC-2) and 3.99 bpw (AC-6) to compare.
@@ 4.2. Metrics
(5) Human Evaluation. For the human evaluation, we focus on evaluating three key aspects of the generated texts: fluency, coherence, and relevance. Our human evaluation was conducted by three researchers who were not involved in the study and each one holds a Bachelor’s degree, demonstrating sufficient expertise in the NLP field. The evaluation process involved a mixed assessment sheet containing both generated stego and cover texts, only stego text scores were selected for analysis. The evaluators, unaware of the underlying algorithms, labels, or any other details, were tasked solely with assessing and scoring the text quality. Each annotator independently rates the texts on a five-point scale (ranging from ’very poor’ to ’very good’) for fluency, coherence, and relevance. The higher evaluation score denotes the better generated stego texts.
@@ 4.3. Results and Analysis
[…]
Table 1 demonstrates the results of text quality of the generated stego texts. In the exploration of literature (Yang et al., 2020b), it is found that texts on public social networks are written by people of different ages and backgrounds in different ways of expression. This leads to the fact that most of the human-written sentences may not obey the optimal language model and form a large variance. From the security, if the generated stego texts have more similar PPL values with normal sentences, their security is higher.
In addition, the higher similarity represents the more similarity between cover and stego sentences, namely, higher security. In our experiments, the average PPL of the normal sentence in the News dataset is 185.64. The “AC-2” can achieve the most similar PPL value and the lowest SS. In addition, the “Discop” algorithm gains the best SS and the maximum gap of PPL. Our method underperforms the Discop because Discop uses the same prefix but different random numbers to generate cover and stego texts, resulting in similar semantics. In contrast, our method generates stego and cover texts using different keywords, thereby leading to slightly lower similarity. Additional experiments confirmed that our method achieves a semantic similarity score of 0.9103 when using the same keywords, which is better than Discop. While the proposed LLM-Stega can make the best trade-off between PPL and SS.
Table 3. Comparison of Steganographic Examples Generated by Proposed Method and Existing Methods.
Steganography Stego-text example
AC in the past few years of the past year and is the most comprehensive issue of a generation of the
ADG j d vance a man who tells us about the constitution stresses there is no indication of collapse
Discop The play begins with having a small boy entering a bar in a night club. Out of the three bars he takes, he is attracted to one woman. Nils André, one of the dancers, is attracted to him. ”Nils is as nice as I am,” André says. The trio sing and dance together in a night club as the bar explodes. der Garçons, an act of urban youth, also appear in the bar as well as at a church. This is where Nils tells the woman
[…]
@@
(4) Human Evaluation. In order to further evaluate the effectiveness of the proposed LLM-Stega, we implement the human evaluation experiment. The details of evaluation scores are illustrated in Figure. 4. The experimental results indicate that the LLM-Stega outperformed the other three algorithms. This suggests that LLM-Stega is more effective in generating steganographic texts that are contextually relevant and less detectable. The results demonstrate LLM-Stega’s superior performance in generating steganographic texts that are fluent, coherent, and relevant. This finding is significant in the context of text steganography, highlighting the potential of LLM-Stega in applications where the imperceptibility of the embedded message is crucial. Besides, in Table 3, we also randomly select some examples of stego text generated by LLM-Stega and existing methods for qualitative analysis. We find that the stego text is fluent enough, with correct grammar and coherent semantics.
@@ 4.4. Ablation Study
The experimental results are shown in Table LABEL:tab:AIFM. We notice that using ”Initial optimization” increases the PPL value and decreases BERT scores, namely, the text quality is reduced. This is because the main goal of “Initial optimization” is to ensure accurately extracting secret messages. In subsequent optimization phases, we aim to improve text quality on the condition that the secret information is extracted accurately, achieving high imperceptibility for the stego-text. Thus, both the PPL and the reject rate have decreased. Besides, the single instance of rejection sampling rate is due to the tuned prompt’s ability to produce text that allows for 100% extraction of the secret information in just one attempt. The experimental results show that the “further optimization” process meets expectations. After ”deep iterations”, the probability of rejection sampling drops to its lowest, while the quality of the text is enhanced to its highest. The experimental results demonstrated that the proposed prompt optimization based on reject sampling significantly improved the quality of the generated stego texts.
@@ Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China under Grant 62272463.